lori is concidering 2 different cellular telephone plans.
plan (a) has a monthly charge of $19.95 plus 33 cents per minute.
plan (b) has a monthly charge of $29.95 plus 10 cents per minute. how long would lori have to talk in a month for the two plans to have the same total cost?

Answer by rfer(16322)

19.95+0.33x=29.95+0.10x
0.23x=10
x=10/0.23
x=43.5 minutes